# VisMem: Latent Vision Memory Unlocks Potential of Vision-Language Models

### v1

@@ 8.3 Baselines
We select a total of 16 baselines, including the vanilla model [bai2025qwen2], 5 direct training paradigms: SFT, Visual-RFT [liu2025visual], VLM-R1 [shen2025vlm], Vision-R1 [huang2025vision], and PAPO [wang2025perception]; 5 image-level paradigms: Sketchpad [hu2024visual], GRIT [fan2025grit], PixelReasoner [su2025pixel], DeepEyes [zheng2025deepeyes], and OpenThinkImg [su2025openthinkimg]; 4 token-level paradigms: Scaffold [lei2024scaffolding], ICoT [gao2025interleaved], MINT-CoT [chen2025mint], and VPT [yu2025introducing]; and 1 latent space paradigm: Mirage [yang2025machine].
Here, VLM-R1 [shen2025vlm] and Vision-R1 [huang2025vision] follow the main GRPO [guo2025deepseek] paradigm based on VLMs. To assess the effectiveness of different methods, our VisMem is trained on Qwen-2.5-VL-7B [bai2025qwen2]. For strategies initially implemented on other base models, e.g., GPT-4o [hurst2024gpt] and Qwen2-VL [wang2024qwen2], we transfer them to Qwen2.5-VL-7B [bai2025qwen2] for fair comparison. Besides, we maintain identical training datasets across various counterparts; however, for those with curated datasets, we follow their original settings. For example, Mirage [yang2025machine] requires additional labeled training images, so we follow its original training dataset; GRIT [fan2025grit] uses a tailored training process with designed data; and MINT-CoT [chen2025mint] curates high-quality mathematical samples with grids and annotations.
@@ 9.1 Benchmark Subset Results towards Visual Sub-capacities
To precisely identify the capability boundaries and advantages of our VisMem, rather than relying solely on overall scores to judge its quality, we evaluate the results of subsets of MuirBench [wang2025muirbench] and LogicVista [xiao2024logicvista] benchmarks. We select 9 subsets of the former benchmark, including: counting, grounding, matching, scene, difference, cartoon, diagram, geographic, and retrieval. While in the latter benchmark, we also select 10 subsets, including 5 reasoning skills: inductive, deductive, numerical, spatial, and mechanical, and 5 capacities: patterns, puzzles, OCR, graphs, and tables. It is worth noting that the selected subsets are only part of the benchmark, thus, the average values of the 10 subsets are not the results of the benchmarks.
As listed in Tab. 5, compared with VLM-R1 [shen2025vlm] and Vision-R1 [huang2025vision], our VisMem achieves the best results on 7 subsets and ranks second on the remaining two subsets. Specifically, it has a generalized enhancement of at least 5% over the base model. Besides, VisMem improves the performance the vanilla model by 16.7% / 18.2% / 11.8% / 13.7% on the counting, grounding, geographic, and retrieval sub-tasks, vastly exceeding the second-best counterpart by 7.0-13.1%. These results indicate that our latent vision memory system significantly promote the fine-grained visual cognition and perception of the base VLMs.
As presented in Tab. 6, our VisMem outperforms two baseline models, i.e., VLM-R1 [shen2025vlm] and Vision-R1 [huang2025vision], by achieving the top performance across 8 subsets. Specifically, it delivers a generalized improvement of no less than 7% over the base model. Notably, on inductive, deductive, graph-based, and table-based sub-tasks, VisMem surpasses the vanilla model by 14.8%, 14.8%, 18.4%, and 21.1%, respectively, which exceeds the second-ranked model by a substantial margin of 5.3–7.1%. These results demonstrate that our latent visual memory system delivers contextualized semantic knowledge, thereby enhancing visual reasoning and robust generation capabilities.
Table 7: Results of various models with full training datasets and partial datasets (Visual CoT [shao2024visual] and Mulberry [yao2024mulberry]), and evaluated across four benchmarks.
@@
Visual-RFT [liu2025visual] 70.5 65.3 62.9 57.8 26.5 24.2 70.7 66.0
VLM-R1 [shen2025vlm] 73.0 67.7 63.8 59.0 34.6 32.1 69.9 66.1
Vision-R1 [huang2025vision] 71.7 68.4 64.0 59.8 38.7 35.6 72.6 67.1
PAPO [wang2025perception] 69.8 68.6 56.7 56.4 34.8 32.8 67.7 66.4
DeepEyes [zheng2025deepeyes] 70.5 67.9 63.0 60.6 31.5 27.9 72.6 68.5
OpenThinkImg [su2025openthinkimg] 71.6 69.5 61.7 59.7 28.0 25.9 74.0 68.3
ICoT [gao2025interleaved] 67.9 67.1 57.0 56.4 30.8 28.3 69.1 68.4
MINT-CoT [chen2025mint] 69.5 68.4 58.9 57.8 39.2 36.4 71.4 70.2
Mirage [yang2025machine] 71.8 70.2 59.0 57.2 35.4 33.1 66.1 64.0
@@
Table 8: Results of various models on MMVet [yu2024mmvet] with four-stage continual learning. Stage 0: MMVet [yu2024mmvet]; Stage 1: BLINK [fu2024blink], and MuirBench [wang2025muirbench]; Stage 2: LogicVista [xiao2024logicvista], and Math-V [wang2024measuring]; Stage 3: MultiTrust [zhang2024multitrust], and MMVU [liu2025unveiling].
@@
Visual-RFT [liu2025visual] 74.0 72.2 67.3 65.7 70.5
VLM-R1 [shen2025vlm] 77.8 74.1 66.4 66.9 73.0
Vision-R1 [huang2025vision] 76.9 74.0 66.1 66.3 71.7
PAPO [wang2025perception] 75.0 74.5 63.4 62.9 69.8
DeepEyes [zheng2025deepeyes] 74.1 74.6 68.9 68.4 70.5
OpenThinkImg [su2025openthinkimg] 76.2 74.7 66.5 67.9 71.6
ICoT [gao2025interleaved] 71.9 71.3 67.1 64.7 67.9
MINT-CoT [chen2025mint] 72.4 71.8 65.8 66.2 69.5
Mirage [yang2025machine] 79.1 77.8 68.7 67.0 71.8
@@
Figure 7: Results of various models of the cross-domain generalization study. Models are only trained on Visual CoT [shao2024visual] and Mulberry [yao2024mulberry], and are evaluated on four benchmarks.
Figure 8: Results of four-stage continual learning on MMVet [yu2024mmvet]. The model is sequentially trained on each training data combination (Stage 0 $\to$ Stage 1 $\to$ Stage 2 $\to$ Stage 3). Stage 0 only includes MMVet [yu2024mmvet] as training data, while Stage 1, 2, 3 add data targeting visual understanding [chen2024we, ying2024mmt, fu2024blink, wang2025muirbench], reasoning [yue2024mmmu, xiao2024logicvista, wang2024measuring, wang2025mv],generation [guan2024hallusionbench, zhang2024multitrust, liu2025unveiling].
@@ 9.2 Cross-domain Generalization
To evaluate the cross-domain generalization capability of our model, we train it exclusively on general datasets, namely, Visual CoT [shao2024visual] and Mullberry [yao2024mulberry]), to verify whether latent visual memory can be transferred to unseen domains. As shown in Tab. 7 and Fig. 8, our method demonstrates superior performance, which exhibits a smaller performance drop than the fully trained model across all four selected benchmarks, confirming strong cross-domain generalization. Despite being trained on only two datasets, our method achieves a significant performance improvement of 9.1–20.5% across the four benchmarks, with a mere 2% performance gap relative to the fully trained model. When compared to other baselines, it still maintains a performance lead of 3.4% / 6.7% / 2.7% / 4.7% across the four evaluations, respectively.
In general, the image-level, token-level, and latent space paradigms suffer from smaller performance degradation, whereas the direct training paradigm exhibits inferior generalization ability. For example, VLM-R1 [shen2025vlm] experiences a 5.3% performance drop; by contrast, this value is only 2.1% for OpenThinkImg [su2025openthinkimg], 1.1% for MINT-CoT [chen2025mint], and 2.3% for our method. These results indicate that while direct training optimizations notably improve performance on specific tasks, they compromise generalization ability to some extent.
@@ 9.3 Catastrophic Forgetting Mitigation
To assess the extent of catastrophic forgetting, we conducted continual learning experiments with our VisMem and other baselines. As presented in Tab. 8 and Fig. 8, our method effectively mitigates forgetting of earlier tasks. It consistently achieves the best performance at each stage, demonstrating strong robustness against catastrophic forgetting. Following four-stage sequential continual training, it retains 72.1% performance on MMVet [yu2024mmvet], outperforming 68.4% of DeepEyes [zheng2025deepeyes] and 67.0% of Mirage [yang2025machine].
While the direct training paradigm significantly improves performance on specific tasks, it adapts to new tasks via direct updates to core parameters. This introduces conflicts when parameter update directions contradict the storage of existing knowledge, compounded by a lack of constraints from prior knowledge. Consequently, in stage 3, the performance of most direct training methods even falls below that of the vanilla model. In contrast, methods such as OpenThinkImg [su2025openthinkimg] and our proposed VisMem exhibit stronger knowledge retention and forward transfer capabilities. For instance, in stage 3, training on additional datasets further improves their performance on MMVet [yu2024mmvet].
@@ 9.4 Versatility across Various Base Models
As presented in Tab. 2 and Fig. 11, we incorporate our latent visual memory paradigm into 9 base models, including Qwen2.5-VL-3B/7B/32B [bai2025qwen2], LLaVA-OV-1.5-4B/8B [an2025llava], and InternVL-3.5-4B/8B/14B/38B [wang2025internvl3]. Our VisMem consistently enhances the visual capabilities of all base models, spanning 3B to 38B parameter sizes across three VLM families. For the widely used medium-sized models (i.e., 7B or 8B parameter models), our latent visual memory delivers substantial performance gains, which brings a 6.3–23.1% improvement across all benchmarks for Qwen2.5-VL-7B [bai2025qwen2], a 5.5–20.2% improvement for LLaVA-OV-1.5-8B [an2025llava], and a 4.8–17.6% improvement for InternVL-3.5-8B [wang2025internvl3], respectively.
@@
Table 9: Ablations of latent vision memory invocation and dual vision memory formation. Following [zhang2025memgen], “Random Invocation” denotes that the latent memory is inserted into the output sequence with a certain probability when outputting delimiter symbol tokens, and short or long latent memory is inserted with equal probability. When only utilizing short or long latent memory, we directly skip the formation of the specific memory if invocation tokens are predicted and continue the process of decoding.
@@ 9.5 Ablation Study
The vanilla model establishes a baseline characterized by the shortest inference time and highest speed across all benchmarks, yet exhibits the lowest performance. This confirms that latent vision memory is indispensable for enhancing task performance. For the random memory invocation variants, increasing the invocation probability (25%–100%) results in longer inference time and reduced speed. Performance peaks at a 75% probability before declining, indicating that excessive memory invocation impairs efficiency without yielding additional performance benefits. Ablation studies of the short-term and long-term memory components reveal task-specific advantages: the short-term memory component outperforms on MuirBench [wang2025muirbench] and MultiTrust [zhang2024multitrust], while the long-term component demonstrates superior performance on MV-Math [wang2025mv]. Notably, the complete VisMem framework achieves the highest performance across all benchmarks, validating the value of integrating dual-component vision memory for balanced and robust visual capacities.
@@ 9.7 Sensitive Analysis of Sequence Lengths
We conduct an analysis on MMVet [yu2024mmvet] focused on the lengths of three key sequences: the memory query $K$, the short-term latent visual memory $N_{s}$, and the long-term latent visual memory $N_{l}$. It is observed that as the lengths of these three sequences increase from 2 to 32, model performance improves accordingly, but this is accompanied by increased computational costs.
@@ 9.8 Inference Efficiency
Visual-RFT [liu2025visual] 0.76 1.32 70.5 3.81 0.26 62.9 5.66 0.17 26.5 3.65 0.27 70.7
VLM-R1 [shen2025vlm] 0.77 1.30 73.0 3.83 0.26 63.8 7.88 0.13 34.6 3.69 0.27 69.9
Vision-R1 [huang2025vision] 0.77 1.30 71.7 3.83 0.26 64.0 8.42 0.12 38.7 3.71 0.27 72.6
PAPO [wang2025perception] 0.76 1.32 69.8 3.81 0.26 56.7 6.74 0.15 34.8 3.68 0.27 67.7
Sketchpad [hu2024visual] 2.39 0.42 64.5 8.90 0.11 52.8 9.10 0.11 24.6 5.47 0.18 66.2
GRIT [fan2025grit] 0.80 1.25 67.8 4.07 0.25 51.0 8.45 0.12 22.4 4.06 0.25 67.3
PixelReasoner [su2025pixel] 1.45 0.69 67.1 7.34 0.14 60.5 9.96 0.10 25.9 5.60 0.18 69.9
DeepEyes [zheng2025deepeyes] 3.21 0.31 70.5 8.46 0.12 63.0 11.72 0.09 31.5 6.14 0.16 72.6
OpenThinkImg [su2025openthinkimg] 3.68 0.27 71.6 8.69 0.12 61.7 10.38 0.10 28.0 6.43 0.16 74.0
Scaffold [lei2024scaffolding] 0.83 1.20 67.0 4.35 0.23 52.9 7.01 0.14 21.0 3.88 0.26 68.5
ICoT [gao2025interleaved] 0.97 1.15 67.9 4.57 0.22 57.0 8.94 0.11 30.8 4.20 0.24 69.1
MINT-CoT [chen2025mint] 0.81 1.23 69.5 4.18 0.24 58.9 7.89 0.13 39.2 4.03 0.25 71.4
VPT [yu2025introducing] 2.98 0.34 70.8 9.63 0.10 63.5 9.59 0.10 34.7 5.79 0.17 64.7
Mirage [yang2025machine] 0.86 1.16 71.8 4.02 0.25 59.0 7.71 0.13 35.4 3.82 0.26 66.1

### v2

@@ 8.3 Baselines
We select a total of 16 baselines, including the vanilla model [4], 5 direct training paradigms: SFT, Visual-RFT [35], VLM-R1 [44], Vision-R1 [26], and PAPO [66]; 5 image-level paradigms: Sketchpad [24], GRIT [13], PixelReasoner [48], DeepEyes [87], and OpenThinkImg [49]; 4 token-level paradigms: Scaffold [28], ICoT [16], MINT-CoT [9], and VPT [75]; and 1 latent space paradigm: Mirage [70].
Here, VLM-R1 [44] and Vision-R1 [26] follow the main GRPO [20] paradigm based on VLMs. To assess the effectiveness of different methods, our VisMem is trained on Qwen-2.5-VL-7B [4]. For strategies initially implemented on other base models, e.g., GPT-4o [27] and Qwen2-VL [61], we transfer them to Qwen2.5-VL-7B [4] for fair comparison. Besides, we maintain identical training datasets across most counterparts; however, for those three methods with specially curated datasets, we follow their original settings. Namely, Mirage [70] requires additional labeled training images, so we follow its original training dataset; GRIT [13] uses a tailored training process with designed data; and MINT-CoT [9] curates high-quality mathematical samples with grids and annotations.
@@ 9.1 Benchmark Subset Results towards Visual Sub-capacities
To precisely identify the capability boundaries and advantages of our VisMem, rather than relying solely on overall scores to judge its quality, we evaluate the results of subsets of MuirBench [57] and LogicVista [67] benchmarks. We select 9 subsets of the former benchmark, including: counting, grounding, matching, scene, difference, cartoon, diagram, geographic, and retrieval. While in the latter benchmark, we also select 10 subsets, including 5 reasoning skills: inductive, deductive, numerical, spatial, and mechanical, and 5 capacities: patterns, puzzles, OCR, graphs, and tables. It is worth noting that the selected subsets are only part of the benchmark, thus, the average values of the 10 subsets are not the results of the benchmarks.
As listed in Tab. 5, compared with VLM-R1 [44] and Vision-R1 [26], our VisMem achieves the best results on 7 subsets and ranks second on the remaining two subsets. Specifically, it has a generalized enhancement of at least 5% over the base model. Besides, VisMem improves the performance the vanilla model by 16.7% / 18.2% / 11.8% / 13.7% on the counting, grounding, geographic, and retrieval sub-tasks, vastly exceeding the second-best counterpart by 7.0-13.1%. These results indicate that our latent vision memory system significantly promote the fine-grained visual cognition and perception of the base VLMs.
As presented in Tab. 6, our VisMem outperforms two baseline models, i.e., VLM-R1 [44] and Vision-R1 [26], by achieving the top performance across 8 subsets. Specifically, it delivers a generalized improvement of no less than 7% over the base model. Notably, on inductive, deductive, graph-based, and table-based sub-tasks, VisMem surpasses the vanilla model by 14.8%, 14.8%, 18.4%, and 21.1%, respectively, which exceeds the second-ranked model by a substantial margin of 5.3–7.1%. These results demonstrate that our latent visual memory system delivers contextualized semantic knowledge, thereby enhancing visual reasoning and robust generation capabilities.
Table 7: Results of various models with full training datasets and partial datasets (Visual CoT [42] and Mulberry [71]), and evaluated across four benchmarks.
@@
Visual-RFT [35] 70.5 65.3 62.9 57.8 26.5 24.2 70.7 66.0
VLM-R1 [44] 73.0 67.7 63.8 59.0 34.6 32.1 69.9 66.1
Vision-R1 [26] 71.7 68.4 64.0 59.8 38.7 35.6 72.6 67.1
PAPO [66] 69.8 68.6 56.7 56.4 34.8 32.8 67.7 66.4
DeepEyes [87] 70.5 67.9 63.0 60.6 31.5 27.9 72.6 68.5
OpenThinkImg [49] 71.6 69.5 61.7 59.7 28.0 25.9 74.0 68.3
ICoT [16] 67.9 67.1 57.0 56.4 30.8 28.3 69.1 68.4
MINT-CoT [9] 69.5 68.4 58.9 57.8 39.2 36.4 71.4 70.2
Mirage [70] 71.8 70.2 59.0 57.2 35.4 33.1 66.1 64.0
@@
Table 8: Results of various models on MMVet [76] with four-stage continual learning. Stage 0: MMVet [76]; Stage 1: BLINK [15], and MuirBench [57]; Stage 2: LogicVista [67], and Math-V [59]; Stage 3: MultiTrust [82], and MMVU [34].
@@
Visual-RFT [35] 74.0 72.2 67.3 65.7 70.5
VLM-R1 [44] 77.8 74.1 66.4 66.9 73.0
Vision-R1 [26] 76.9 74.0 66.1 66.3 71.7
PAPO [66] 75.0 74.5 63.4 62.9 69.8
DeepEyes [87] 74.1 74.6 68.9 68.4 70.5
OpenThinkImg [49] 76.2 74.7 66.5 67.9 71.6
ICoT [16] 71.9 71.3 67.1 64.7 67.9
MINT-CoT [9] 72.4 71.8 65.8 66.2 69.5
Mirage [70] 79.1 77.8 68.7 67.0 71.8
@@
Figure 7: Results of various models of the cross-domain generalization study. Models are only trained on Visual CoT [42] and Mulberry [71], and are evaluated on four benchmarks.
Figure 8: Results of four-stage continual learning on MMVet [76]. The model is sequentially trained on each training data combination (Stage 0 $\to$ Stage 1 $\to$ Stage 2 $\to$ Stage 3). Stage 0 only includes MMVet [76] as training data, while Stage 1, 2, 3 add data targeting visual understanding [15, 57], reasoning [67, 59], and generation [82, 34].
@@ 9.2 Cross-domain Generalization
To evaluate the cross-domain generalization capability of our model, we train it exclusively on general datasets, namely, Visual CoT [42] and Mullberry [71]), to verify whether latent visual memory can be transferred to unseen domains. As shown in Tab. 7 and Fig. 8, our method demonstrates superior performance, which exhibits a smaller performance drop than the fully trained model across all four selected benchmarks, confirming strong cross-domain generalization. Despite being trained on only two datasets, our method achieves a significant performance improvement of 9.1–20.5% across the four benchmarks, with a mere 2% performance gap relative to the fully trained model. When compared to other baselines, it still maintains a performance lead of 3.4% / 6.7% / 2.7% / 4.7% across the four evaluations, respectively.
In general, the image-level, token-level, and latent space paradigms suffer from smaller performance degradation, whereas the direct training paradigm exhibits inferior generalization ability. For example, VLM-R1 [44] experiences a 5.3% performance drop; by contrast, this value is only 2.1% for OpenThinkImg [49], 1.1% for MINT-CoT [9], and 2.3% for our method. These results indicate that while direct training optimizations notably improve performance on specific tasks, they compromise generalization ability to some extent.
@@ 9.3 Catastrophic Forgetting Mitigation
To assess the extent of catastrophic forgetting, we conducted continual learning experiments with our VisMem and other baselines. As presented in Tab. 8 and Fig. 8, our method effectively mitigates forgetting of earlier tasks. It consistently achieves the best performance at each stage, demonstrating strong robustness against catastrophic forgetting. Following four-stage sequential continual training, it retains 72.1% performance on MMVet [76], outperforming 68.4% of DeepEyes [87] and 67.0% of Mirage [70].
While the direct training paradigm significantly improves performance on specific tasks, it adapts to new tasks via direct updates to core parameters. This introduces conflicts when parameter update directions contradict the storage of existing knowledge, compounded by a lack of constraints from prior knowledge. Consequently, in stage 3, the performance of most direct training methods even falls below that of the vanilla model. In contrast, methods such as OpenThinkImg [49] and our proposed VisMem exhibit stronger knowledge retention and forward transfer capabilities. For instance, in stage 3, training on additional datasets further improves their performance on MMVet [76].
@@ 9.4 Versatility across Various Base Models
As presented in Tab. 2 and Fig. 11, we incorporate our latent visual memory paradigm into 9 base models, including Qwen2.5-VL-3B/7B/32B [4], LLaVA-OV-1.5-4B/8B [1], and InternVL-3.5-4B/8B/14B/38B [63]. Our VisMem consistently enhances the visual capabilities of all base models, spanning 3B to 38B parameter sizes across three VLM families. For the widely used medium-sized models (i.e., 7B or 8B parameter models), our latent visual memory delivers substantial performance gains, which brings a 6.3–23.1% improvement across all benchmarks for Qwen2.5-VL-7B [4], a 5.5–20.2% improvement for LLaVA-OV-1.5-8B [1], and a 4.8–17.6% improvement for InternVL-3.5-8B [63], respectively.
@@
Table 9: Ablations of latent vision memory invocation and dual vision memory formation. Following [81], “Random Invocation” denotes that the latent memory is inserted into the output sequence with a certain probability when outputting delimiter symbol tokens, and short or long latent memory is inserted with equal probability. When only utilizing short or long latent memory, we directly skip the formation of the specific memory if invocation tokens are predicted and continue the process of decoding.
@@ 9.5 Ablation Study
The vanilla model establishes a baseline characterized by the shortest inference time and highest speed across all benchmarks, yet exhibits the lowest performance. This confirms that latent vision memory is indispensable for enhancing task performance. For the random memory invocation variants, increasing the invocation probability (25%–100%) results in longer inference time and reduced speed. Performance peaks at a 75% probability before declining, indicating that excessive memory invocation impairs efficiency without yielding additional performance benefits. Ablation studies of the short-term and long-term memory components reveal task-specific advantages: the short-term memory component outperforms on MuirBench [57] and MultiTrust [82], while the long-term component demonstrates superior performance on MV-Math [60]. Notably, the complete VisMem framework achieves the highest performance across all benchmarks, validating the value of integrating dual-component vision memory for balanced and robust visual capacities.
@@ 9.7 Sensitive Analysis of Sequence Lengths
We conduct an analysis on MMVet [76] focused on the lengths of three key sequences: the memory query $K$, the short-term latent visual memory $N_{s}$, and the long-term latent visual memory $N_{l}$. It is observed that as the lengths of these three sequences increase from 2 to 32, model performance improves accordingly, but this is accompanied by increased computational costs.
@@ 9.8 Inference Efficiency
Visual-RFT [35] 0.76 1.32 70.5 3.81 0.26 62.9 5.66 0.17 26.5 3.65 0.27 70.7
VLM-R1 [44] 0.77 1.30 73.0 3.83 0.26 63.8 7.88 0.13 34.6 3.69 0.27 69.9
Vision-R1 [26] 0.77 1.30 71.7 3.83 0.26 64.0 8.42 0.12 38.7 3.71 0.27 72.6
PAPO [66] 0.76 1.32 69.8 3.81 0.26 56.7 6.74 0.15 34.8 3.68 0.27 67.7
Sketchpad [24] 2.39 0.42 64.5 8.90 0.11 52.8 9.10 0.11 24.6 5.47 0.18 66.2
GRIT [13] 0.80 1.25 67.8 4.07 0.25 51.0 8.45 0.12 22.4 4.06 0.25 67.3
PixelReasoner [48] 1.45 0.69 67.1 7.34 0.14 60.5 9.96 0.10 25.9 5.60 0.18 69.9
DeepEyes [87] 3.21 0.31 70.5 8.46 0.12 63.0 11.72 0.09 31.5 6.14 0.16 72.6
OpenThinkImg [49] 3.68 0.27 71.6 8.69 0.12 61.7 10.38 0.10 28.0 6.43 0.16 74.0
Scaffold [28] 0.83 1.20 67.0 4.35 0.23 52.9 7.01 0.14 21.0 3.88 0.26 68.5
ICoT [16] 0.97 1.15 67.9 4.57 0.22 57.0 8.94 0.11 30.8 4.20 0.24 69.1
MINT-CoT [9] 0.81 1.23 69.5 4.18 0.24 58.9 7.89 0.13 39.2 4.03 0.25 71.4
VPT [75] 2.98 0.34 70.8 9.63 0.10 63.5 9.59 0.10 34.7 5.79 0.17 64.7
Mirage [70] 0.86 1.16 71.8 4.02 0.25 59.0 7.71 0.13 35.4 3.82 0.26 66.1